# Prospects for Probing Sub-GeV Leptophilic Dark Matter with the Future VLAST

### v1

@@ I Introduction
The search for non-gravitational signals from dark matter (DM) is essential to understanding its fundamental properties. Recent efforts on hunting Weakly Interacting Massive Particles (WIMPs) in the mass between $\,\mathrm{GeV}$ to $\,\mathrm{TeV}$ range Chang et al. (2008); Aharonian et al. (2009); Adriani et al. (2009); Hooper and Goodenough (2011); Zhou et al. (2015); Calore et al. (2015); Daylan et al. (2016); Cui et al. (2017); Cuoco et al. (2017); Cholis et al. (2019); Fan et al. (2022, 2024) have uncovered intriguing anomalies, though none have definitively resolved the DM puzzle. In parallel, advances in detector technology have greatly improved sensitivity to new physics below the electroweak scale, generating growing interest in light DM. Among the viable paradigm, thermal relic sub-GeV DM is especially compelling, as it naturally emerges from early-universe thermodynamics Boddy and Kumar (2015); Bartels et al. (2017); Matsumoto et al. (2019); Chen et al. (2018); Cirelli et al. (2021); Coogan et al. (2021); Caputo et al. (2023); Berlin et al. (2024); De la Torre Luque et al. (2024); Chen et al. (2024); O’Donnell and Slatyer (2025); Balan et al. (2025); Saha (2025); Cirelli and Kar (2025); Watanabe et al. (2025).
@@
Direct detection of sub-GeV DM remains challenging due to the suppressed recoil energies from DM-nucleon interactions, which fall below the thresholds of current detectors. As a result, indirect detection via annihilation or decay products has gained attention as a complementary strategy. As early as the 1990s, instruments like COMPTEL Strong et al. (1999) and EGRET Hunter et al. (1997) conducted gamma-ray observations in the range of MeV to GeV. However, their sensitivities were several orders of magnitude lower than those of higher-energy experiments such as Fermi-LAT Atwood et al. (2009) and DAMPE Chang et al. (2017), leaving an observational gap known as the “MeV Gap”. In recent years, several next-generation gamma-ray telescopes have been proposed to close this gap and enhance sensitivity to MeV-scale gamma rays, including COSI Tomsick (2021), the Very Large Area gamma-ray Space Telescope (VLAST) Fan et al. (2022); Pan et al. (2024), e-ASTROGAM Tavani et al. (2018), GECCO Orlando et al. (2022), AMEGO Kierans (2020), GRAMS Aramaki et al. (2020), AdEPT Hunter et al. (2014), and MAST Dzhatdoev and Podlesnyi (2019).
@@
VLAST Fan et al. (2022); Pan et al. (2024) is a space mission designed to detect gamma-ray photons through both Compton scattering and electron-positron pair production mechanisms. The mission aims to perform a comprehensive all-sky survey from low-Earth orbit using three primary detector components: an anti-coincidence detector to suppress charged-particle background, a Silicon tracker coupled with a low-energy gamma-ray detector for accurate reconstruction of particle trajectories and energy measurement at low energies, and a high-energy imaging calorimeter to measure energy deposition and reconstruct the electromagnetic showers induced by high-energy particles. VLAST builds upon the foundational principles and structure of previous gamma-ray detector but introduces a key modification which replacing the tungsten converter in the tracker with thin CsI tiles. This modification enable energy measurement of Compton electrons, along with simultaneous both energy and direction measurements of Compton photons, thereby allowing precise reconstruction of Compton events, especially for the photon energy around $\mathcal{O}(\,\mathrm{MeV})$. This detector configuration enable simultaneous observation across a broad energy range from MeV to TeV, supporting the key scientific objectives such as searching for DM signatures, monitoring of transient and persistent astrophysical phenomena, probing cosmic ray origins and transportation, studying cosmic evolution, and testing fundamental physical laws.
In the early universe, charged particles produced by DM annihilation can interact with the cosmic microwave background (CMB) photons, distorting its black-body spectrum Chluba and Sunyaev (2012). Such energy injection during recombination also modifies the thermal and ionization history of the primordial plasma, leaving detectable signatures in the CMB temperature and polarization anisotropies Adams et al. (1998); Chen and Kamionkowski (2004); Padmanabhan and Finkbeiner (2005); Slatyer et al. (2009). For the standard thermal paradigm, the typical DM annihilation cross section is $\langle\sigma v\rangle\sim 10^{-26}~{}\mathrm{cm}^{3}/\mathrm{s}$ Feng and Kumar (2008); Arcadi et al. (2018); Cirelli et al. (2024), whereas the angular power spectrum measurements of Planck collaboration constrain the $s$-wave annihilation cross section to $\langle\sigma v\rangle\times(10~{}\mathrm{GeV}/m_{\mathrm{DM}})\lesssim 3\times 10^{-26}~{}\mathrm{cm}^{3}/\mathrm{s}$ for the $e^{+}e^{-}$ channel Aghanim et al. (2020a). These measurements place stringent bounds on the annihilation cross section of sub-GeV DM Slatyer et al. (2009); Slatyer (2016); Steigman (2015); Aghanim et al. (2020b); Cang et al. (2020); Kawasaki et al. (2021); Matsumoto et al. (2023); Chen et al. (2024).
@@
In this work, we investigate the sensitivity of VLAST to sub-GeV leptophilic DM annihilation in DM-rich dwarf galaxy, such as the Draco. We consider three representative DM interactions: (i) scalar DM with a scalar mediator, (ii) Dirac DM with a scalar mediator, and (iii) Dirac DM with a vector mediator. These interactions exhibit distinct annihilation behaviors: $s$-wave, $p$-wave, and mixed ($s+p$)-wave processes, respectively. A comprehensive Markov Chain Monte Carlo (MCMC) analysis is performed, incorporating constraints from DM relic abundance, CMB observations, Big Bang Nucleosynthesis (BBN), DM self-interactions, direct detection experiments, and beam dump experiments. Our analysis shows that direct detection experiments like XENON1T impose stringent limits on DM-electron interactions, particularly in the electrophilic scenario. VLAST, however, holds strong potential to probe large regions of the $p$-wave and $(s+p)$-wave annihilation parameter space. Surprisingly, the resonance region offers a promising goal for VLAST. In contrast, the $s$-wave scenario remains challenging due to stringent constraints from CMB observations and relic density requirements.
@@
The remainder of this paper is organized as follows. In Sec. II, we briefly present the key performance characteristics of VLAST and estimate its sensitivity to sub-GeV DM. In Sec. III, we provide an overview of the leptophilic DM Lagrangians considered in this work and outline the constraints incorporated into our likelihood analysis. The allowed parameter space resulting from the likelihood analysis is presented in Sec. IV. Finally, we summarize our conclusions in Sec. VI.
@@ II The VLAST sensitivity prospects
The VLAST detector configuration has been proposed Fan et al. (2022); Pan et al. (2024), and its performance has been evaluated through Monte Carlo simulations using the GEANT4 toolkit, confirming the feasibility of the instrument. VLAST provides an acceptance greater than 10 m2 sr (four times that of Fermi-LAT) and achieves energy and angular resolutions of better than 2% and 0.2 at 10 GeV, respectively. These features significantly enhance its sensitivity to sub-GeV leptophilic dark matter annihilation in DM-rich systems, such as the Draco dwarf galaxy.
The inner Galactic diffuse emission observed by COMPTEL in the $1~{}\,\mathrm{MeV}$ to $30~{}\,\mathrm{MeV}$ range Bouchet et al. (2011) has been analyzed by cross-matching the Swift-BAT hard X-ray catalog with the Fermi-LAT GeV gamma-ray catalog Tsuji et al. (2021). This study includes both matched and unmatched sources, as well as contributions from the Galactic diffuse emission (GDE) and the cosmic gamma-ray background (CGB) Tsuji et al. (2023). The GDE, composed of inverse Compton scattering, bremsstrahlung, and pion decay, is modeled with GALPROP based on Ref. Ackermann et al. (2012). The MeV CGB measured by COMPTEL Weidenspointner et al. (2000) is well described by a broken power-law,
$I(E)=2.2\times 10^{-4}\left(\frac{E}{3~{}\text{MeV}}\right)^{-\Gamma}~{}(\text{MeV cm}^{2}~{}\text{s sr})^{-1},$ (1)
with spectral indices $\Gamma=3.3$ for $E\leq 3~{}\text{MeV}$ and $\Gamma=2$ for $E>3~{}\text{MeV}$. At GeV energies, the CGB is modeled using Fermi-LAT Model A Ackermann et al. (2015a). All-sky maps of each emission component are available online [^1] https://tsuji703.github.io/MeV-All-Sky/files/allsky/Table_allsky.html.
We add a point source at the location of the Draco dwarf galaxy, modeled with a power-law spectrum of index $-2$, in addition to the GDE and CGB components. Using the effective area and angular resolution of VLAST, we simulate the expected count map for one year of observation. The model is then fitted to extract the test statistic (TS) of the point source, defined as
@@
where $L_{\text{GDE+CGB+PS}}$ and $L_{\text{GDE+CGB}}$ are the likelihoods with and without the point source, respectively. By varying the point source normalization until the TS reaches 25 (corresponding to $5\sigma$ significance), we determine the point source sensitivity of VLAST.
@@
Figure 1: Left panel: The sensitivity of VLAST for one year (red dashed line) to gamma-ray emission from the Draco dwarf galaxy, compared with projected sensitivities of e-ASTROGAM for one year (black dashed line) Tavani et al. (2018) and 6-year Fermi-LAT data (blue dashed line) Ackermann et al. (2015b); Albert et al. (2017). Solid lines show the prompt gamma-ray spectra from DM annihilation into lepton pairs assuming a thermal relic cross section of $\langle\sigma v\rangle=2.2\times 10^{-26}~{}\rm cm^{3}\,s^{-1}$ Steigman et al. (2012) with $\mathrm{DM}+\mathrm{DM}\to e^{+}e^{-}$ for two benchmark masses $m_{\rm DM}=10~{}\,\mathrm{MeV}$ (green) and $800~{}\,\mathrm{MeV}$ (red), and $\mathrm{DM}+\mathrm{DM}\to\mu^{+}\mu^{-}$ for $m_{\rm DM}=800~{}\,\mathrm{MeV}$ (blue). Right panel: Projected sensitivities to the DM annihilation cross section. Red lines correspond to VLAST sensitivities for $e^{+}e^{-}$ (solid) and $\mu^{+}\mu^{-}$ (dash-dotted) channels, while black lines show the corresponding sensitivities for e-ASTROGAM.
@@
The left panel of Fig. 1 shows the projected one-year sensitivity of VLAST (red dashed line) and e-ASTROGAM Tavani et al. (2018) (black dashed line), compared to 6-year Fermi-LAT data Ackermann et al. (2015b); Albert et al. (2017) (blue dashed line), for DM annihilation in the Draco dwarf galaxy. The right panel presents the corresponding upper limits on the annihilation cross-section. Note that VLAST shows a sensitivity gap around $10\,\mathrm{MeV}$, resulting from the transition between its two detection mechanisms: Compton scattering and pair production. This transition causes discontinuities in the effective area and angular resolution near $10\,\mathrm{MeV}$, leading to the observed dip in sensitivity.
@@
To highlight the VLAST capability to detect MeV gamma-ray signals from DM annihilation, we compare theoretical flux predictions from the Draco dwarf galaxy, known for its clean astrophysical background. For the DM density distribution, we adopt Navarro-Frenk-White (NFW) profile,
$\rho_{\rm NFW}(r)=\frac{\rho_{s}}{(r/r_{s})(1+r/r_{s})^{2}},$ (3)
@@
where $\rho_{s}=2.96\,\,\mathrm{GeV}\mathrm{/cm}^{3}$ and $r_{s}=728\,\rm pc$ are the scaled density and scaled radius, respectively. The distance from Earth to the center of the dwarf galaxy is $D=76\,\rm kpc$ Pace and Strigari (2019). Therefore, the differential DM-induced gamma-ray flux is given by
[…]
@@
where $dN_{\gamma}/dE_{\gamma}$ is the photon yield spectrum, obtained using the hazma Coogan et al. (2020). The velocity averaged annihilation cross section $\langle\sigma v_{\rm rel.}(r)\rangle$ depends on the relative velocity distribution of DM particles at radius $r$, given by
[…]
@@
where $v_{\rm d}(r)$ is the velocity dispersion derived from Jeans equation Jin et al. (2005).
In the left panel of Fig. 1, we display the prompt emission from DM annihilation into $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ channels, assuming a representative cross section of $\langle\sigma v\rangle=2.2\times 10^{-26}\,\rm cm^{3}\,s^{-1}$ Steigman et al. (2012) for two benchmark DM masses, $m_{\rm DM}=10\,\,\mathrm{MeV}$ and $800\,\,\mathrm{MeV}$. The resulting spectra are shown as solid lines. Note that the DM-induced gamma-ray flux due to inverse Compton scattering with CMB photons is not depicted, as it lies in the X-ray range, below the threshold of VLAST.

### v2

@@ I Introduction
The search for non-gravitational signals from dark matter (DM) is essential to understanding its fundamental properties, particularly its interactions with Standard Model (SM) particles. Within the thermal relic paradigm, DM masses between the MeV and TeV scales can naturally reproduce the observed abundance through freeze-out dynamics Griest:1989wd; Boehm:2003bt; Boehm:2002yz. This framework has long motivated the canonical WIMP program in the GeV to TeV range, where extensive experimental efforts have been carried out Chang:2008aa; HESS:2009chc; PAMELA:2008gwm; Hooper:2010mq; Zhou:2014lva; Calore:2014xka; Daylan:2014rsa; Cui:2016ppb; Cuoco:2016eej; Cholis:2019ejx; Fan:2022dck; Fan:2024rcr; Feng:2008ya; Arcadi:2017kky; Cirelli:2024ssz. Although intriguing anomalies have been reported, a wide array of direct, indirect, and collider searches has severely constrained large portions of the viable parameter space above the GeV scale.
In contrast, thermal relic DM with masses below a GeV is comparatively less explored but still theoretically well motivated. The very light region, at or below the MeV scale, is strongly constrained by cosmological probes such as Big Bang Nucleosynthesis (BBN) and the Cosmic Microwave Background (CMB), leaving only a narrow mass window. The intermediate regime between the MeV and GeV scales, however, remains open  Adams:1998nr; Chen:2003gz; Padmanabhan:2005es; Slatyer:2009yq; Chluba:2011hw; Slatyer:2015jla; Steigman:2015hda; Planck:2018nkj; Cang:2020exa; Kawasaki:2021etm; Matsumoto:2022ojl, even though existing limits from beam dump experiments, low-threshold direct detection, and CMB observations already exclude scenarios with large couplings to the SM particles.
This situation makes the sub-GeV mass range particularly appealing. It arises naturally within the thermal relic framework, has not yet been excluded by current data, and offers a parameter space that is both finite and testable. With upcoming advances across multiple experimental frontiers, the remaining window of sub-GeV DM can be probed decisively in the near future Boddy:2015efa; Bartels:2017dpb; Matsumoto:2018acr; Chen:2018vkr; Cirelli:2020bpc; Coogan:2021sjs; Caputo:2022dkz; Guo:2023kqt; Berlin:2023qco; DelaTorreLuque:2023olp; Chen:2024njd; ODonnell:2024aaw; Balan:2024cmq; Nguyen:2024kwy; Saha:2025wgg; Cirelli:2025qxx; Watanabe:2025pvc. It thus provides a well-motivated and experimentally timely target for the next generation of DM indirect searches.
@@
Historically, instruments like COMPTEL Strong:1998ck and EGRET Hunter:1997qec conducted gamma-ray observations in the range of $\mathcal{O}(1-20)~\,\mathrm{MeV}$. However, their sensitivities were several orders of magnitude lower than those of higher-energy experiments such as Fermi-LAT Fermi-LAT:2009ihh and DAMPE DAMPE:2017cev, leaving an observational gap known as the “MeV Gap”. In recent years, several next-generation gamma-ray telescopes have been proposed to close this gap and enhance sensitivity to MeV-scale gamma rays, including COSI Tomsick:2021wed, the Very Large Area gamma-ray Space Telescope (VLAST) 2022AcASn..63...27F; Pan:2024adp, e-ASTROGAM e-ASTROGAM:2017pxr, GECCO Orlando:2021get, AMEGO Kierans:2020otl, GRAMS Aramaki:2019bpi, AdEPT Hunter:2013wla, and MAST Dzhatdoev:2019kay.
@@
VLAST 2022AcASn..63...27F; Pan:2024adp is a space mission designed to detect gamma-ray photons through both Compton scattering and electron-positron pair production mechanisms. The mission aims to perform a comprehensive all-sky survey from low-Earth orbit using three primary detector components: an anti-coincidence detector to suppress charged-particle background, a Silicon tracker coupled with a low-energy gamma-ray detector for accurate reconstruction of particle trajectories and energy measurement at low energies, and a high-energy imaging calorimeter to measure energy deposition and reconstruct the electromagnetic showers induced by high-energy particles. VLAST builds upon the foundational principles and structure of previous gamma-ray detector but introduces a key modification which replaces the tungsten converter in the tracker with thin CsI tiles. This modification enables energy measurement of Compton electrons, along with simultaneous energy and direction measurements of Compton photons, thereby allowing precise reconstruction of Compton events, especially for the photon energy around $\mathcal{O}(\,\mathrm{MeV})$. This detector configuration enables simultaneous observation across a broad energy range from MeV to TeV, supporting key scientific objectives such as searching for DM signatures, monitoring of transient and persistent astrophysical phenomena, probing cosmic ray origins and transportation, studying cosmic evolution, and testing fundamental physical laws.
[…]
@@
In this work, we investigate the sensitivity of VLAST to sub-GeV leptophilic DM annihilation in DM-rich dwarf galaxy, such as Draco. Based on the systematic construction and classification of 16 renormalizable effective leptophilic models Abdughani:2021oit, we focus on three representative DM interactions: (i) scalar DM with a scalar mediator, (ii) Dirac DM with a scalar mediator, and (iii) Dirac DM with a vector mediator. These interactions exhibit distinct annihilation behaviors: $s$-wave, $p$-wave, and mixed ($s+p$)-wave processes, respectively. A comprehensive Markov Chain Monte Carlo (MCMC) analysis is performed, incorporating constraints from DM relic abundance, CMB observations, BBN, DM self-interactions, direct detection experiments, and beam dump experiments. Our analysis shows that direct detection experiments like XENON1T impose stringent limits on DM-electron interactions, particularly in the electrophilic scenario. VLAST, however, holds strong potential to probe large regions of the $p$-wave and $(s+p)$-wave annihilation parameter space. Surprisingly, the resonance region offers a promising goal for VLAST. In contrast, the $s$-wave scenario remains challenging due to stringent constraints from CMB observations and relic density requirements.
@@
The remainder of this paper is organized as follows. In Sec. II, we briefly present the key performance characteristics of VLAST and estimate its sensitivity to sub-GeV DM. In Sec. III, we provide an overview of the leptophilic DM Lagrangians considered in this work and outline the constraints incorporated into our likelihood function analysis. The allowed parameter space resulting from the likelihood analysis is presented in Sec. IV. Finally, we summarize our conclusions in Sec. VI.
@@ II The VLAST sensitivity prospects
The VLAST detector configuration has been proposed 2022AcASn..63...27F; Pan:2024adp, and its performance has been evaluated through Monte Carlo simulations using the GEANT4 toolkit, confirming the feasibility of the instrument. VLAST provides an acceptance greater than $10~\mathrm{m}^{2}~\mathrm{sr}$ (four times that of Fermi-LAT) and achieves energy and angular resolutions of better than 2% and 0.2 at 10 GeV, respectively. These features significantly enhance its sensitivity to sub-GeV leptophilic DM annihilation in DM-rich systems, such as Draco dwarf galaxy.
Draco dwarf galaxy has an angular size of $\sim 0.26^{\circ}$ 2016JCAP...02..039M, much smaller than the angular resolution of VLAST in the MeV range ($\sim 5^{\circ}$ at 1 MeV and $\sim 3^{\circ}$ at 10 MeV). Therefore, any DM-induced $\gamma$-ray emission from Draco can be treated as a point source for VLAST. We calculate the instrument sensitivity to a point source at Draco location and translate it into limits on the DM annihilation cross section. Previous $\gamma$-ray dwarf galaxy studies Fermi-LAT:2015att; Fermi-LAT:2016uux have shown that sensitivity can be computed for individual energy bins and is independent of the spectral index. Following their approach, to simulate the expected signal, a point source with a power-law spectrum of the form $d\Phi_{\rm PS}/dE_{\gamma}\propto E_{\gamma}^{-2}$ is used as the baseline model for sensitivity calculations.
The astrophysical background model contains the Galactic Diffuse Emission (GDE) and the Cosmic $\gamma$-ray Background (CGB). The GDE component arises from inverse Compton scattering, bremsstrahlung, and $\pi^{0}$ decay, and is modeled with GALPROP following the method of Ref. Fermi-LAT:2012edv. Similarly, following COMPTEL 2000AIPC..510..467W, the CGB component is modeled as a broken power law
$\frac{d\Phi_{\rm CGB}}{dE_{\gamma}}\left(E_{\gamma}\right)=2.2\times 10^{-4}\left(\frac{E_{\gamma}}{3~\text{MeV}}\right)^{-\Gamma}~(\text{MeV cm}^{2}~\text{s sr})^{-1},$ (1)
with photon indices $\Gamma=3.3$ for $E_{\gamma}\leq 3~\mathrm{MeV}$ and $\Gamma=2.0$ for $E_{\gamma}>3~\mathrm{MeV}$. At GeV energies, we adopt the Fermi-LAT Model A spectrum Fermi-LAT:2014ryh. All-sky maps for each component are obtained from publicly available templates  [^1] https://tsuji703.github.io/MeV-All-Sky/files/allsky/Table_allsky.html.
The simulated sky maps include the GDE, CGB, and the point source. The counts are generated using VLAST energy-dependent effective area and point-spread function. A binned likelihood analysis is performed, where the normalizations of GDE and CGB are treated as nuisance parameters.
The test statistic (TS) is defined as
@@
where $L_{\text{GDE+CGB+PS}}$ and $L_{\text{GDE+CGB}}$ are the likelihoods with and without the point source, respectively. By varying the point source normalization until the TS reaches 25 (corresponding to $5\sigma$ significance), we determine the point-source sensitivity of VLAST for one year of survey observations.
@@
Figure 1: Left panel: The sensitivity of VLAST for one year (red dashed line) to gamma-ray emission from Draco dwarf galaxy, compared with projected sensitivities of e-ASTROGAM for one year (black dashed line) e-ASTROGAM:2017pxr and 6-year Fermi-LAT data (blue dashed line) Fermi-LAT:2015att; Fermi-LAT:2016uux. Solid lines show the prompt gamma-ray spectra from DM annihilation into lepton pairs assuming a thermal relic cross section of $\langle\sigma v\rangle=2.2\times 10^{-26}~\rm cm^{3}\,s^{-1}$ Steigman:2012nb with $\mathrm{DM}+\mathrm{DM}\to e^{+}e^{-}$ for two benchmark masses $m_{\rm DM}=10~\,\mathrm{MeV}$ (green) and $800~\,\mathrm{MeV}$ (red), and $\mathrm{DM}+\mathrm{DM}\to\mu^{+}\mu^{-}$ for $m_{\rm DM}=800~\,\mathrm{MeV}$ (blue). Right panel: Projected sensitivities to the DM annihilation cross section. Red lines correspond to VLAST sensitivities for $e^{+}e^{-}$ (solid) and $\mu^{+}\mu^{-}$ (dash-dotted) channels, while black lines show the corresponding sensitivities for e-ASTROGAM.
@@
The left panel of Fig. 1 shows the projected one-year sensitivity of VLAST (red dashed line) and e-ASTROGAM e-ASTROGAM:2017pxr (black dashed line), compared to 6-year Fermi-LAT data Fermi-LAT:2015att; Fermi-LAT:2016uux (blue dashed line), for DM annihilation in Draco dwarf galaxy. The right panel presents the corresponding upper limits on the annihilation cross section. Note that VLAST shows a sensitivity gap around $10\,\mathrm{MeV}$, resulting from the transition between its two detection mechanisms: Compton scattering and pair production. This transition causes discontinuities in the effective area and angular resolution near $10\,\mathrm{MeV}$, leading to the observed dip in sensitivity.
@@
To highlight the VLAST capability to detect MeV gamma-ray signals from DM annihilation, we compare theoretical flux predictions from Draco dwarf galaxy, known for its clean astrophysical background. For the DM density distribution, we adopt Navarro-Frenk-White (NFW) profile,
$\rho_{\rm NFW}(r)=\frac{\rho_{s}}{(r/r_{s})(1+r/r_{s})^{2}}.$ (3)
@@
For Draco, following Ref. Pace:2018tin, we use the scaled density $\rho_{s}=2.96\,\,\mathrm{GeV}\mathrm{/cm}^{3}$, the scaled radius $r_{s}=728\,\rm pc$, and the distance $D=76\,\rm kpc$ from Earth to its center. Therefore, the differential DM-induced gamma-ray flux is given by
$\frac{d\phi_{\gamma}}{dE_{\gamma}}$ $=\int_{0}^{D}\frac{\rho^{2}_{\rm NFW}(r)r^{2}}{4m_{\chi}^{2}D^{2}}\frac{dN_{\gamma}}{dE_{\gamma}}\langle\sigma v_{\rm rel.}(r)\rangle dr,$ (4)
@@
where $dN_{\gamma}/dE_{\gamma}$ is the photon yield spectrum, obtained using Hazma Coogan:2019qpu. The velocity averaged annihilation cross section $\langle\sigma v_{\rm rel.}(r)\rangle$ depends on the relative velocity distribution of DM particles at radius $r$, given by
$\langle\sigma v_{\rm rel}(r)\rangle=\frac{1}{Norm(r)}\int_{0}^{2v_{\rm esc}}\sigma v_{\rm rel}f_{v_{\rm rel.}}(r,v_{\rm rel.})dv_{\rm rel}.$ (5)
@@
where $v_{\rm d}(r)$ is the velocity dispersion derived from Jeans equation Jin:2003sj.
In the left panel of Fig. 1, we display the prompt emission from DM annihilation into $e^{+}e^{-}$ and $\mu^{+}\mu^{-}$ channels, assuming a representative cross section of $\langle\sigma v\rangle=2.2\times 10^{-26}\,\rm cm^{3}\,s^{-1}$ Steigman:2012nb for two benchmark DM masses, $m_{\rm DM}=10\,\,\mathrm{MeV}$ and $800\,\,\mathrm{MeV}$. The resulting spectra are shown as solid lines. Note that the DM-induced gamma-ray flux due to inverse Compton scattering with CMB photons is not depicted, as it lies in the X-ray range, below the threshold of VLAST.